Find the cube roots of 27(cos 279° + i sin 279°).

Help me please
1
Expert's answer
2014-05-14T03:20:28-0400
Using thr De Moivre's formula we get
(27(cos 279° + i sin 279°))^(1/3)=27^(1/3)(cos 279°/3 + i sin 279°/3)=3(cos 93 + i sin 93)=(approximately) -0.157+3i